OpenStudy (anonymous):

Find the sum of the infinite geometric series, if it exists. 4 - 1 + 1/4 - 1/6 + . . . A. - 1 B. 3 C. 16/5 D. does not exist

OpenStudy (anonymous):

4-1+1/4-1/16+........ a=4,r=-1/4 \[\left| r \right|=\frac{ 1 }{ 4 }<1\] \[S \infty=\frac{ a }{1-r }=\frac{ 4 }{1-\left( \frac{ -1 }{4 } \right) }\] \[S \infty=\frac{ 4 }{ 1+\frac{ 1 }{ 4 } }=\frac{ 4 }{\frac{ 4+1 }{ 4 } }=4\times \frac{ 4 }{ 5 }\] \[S \infty=\frac{ 16 }{5 }\]

OpenStudy (anonymous):

thank you for explaining @surjithayer ;-)